OpenStudy (anonymous):

Evaluate the limit lim x->0 (x+8)^(1/3) -2 ___________ x

OpenStudy (anonymous):

This? \[\lim_{x \rightarrow 0} \frac{(x+8)^{1/3}-2}{x}\]

OpenStudy (anonymous):

Yes i was just going to type that, sorry. I've been reviewing limits and I keep screwing this one up.

OpenStudy (anonymous):

You can use l'hospital's rule

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{(x+8)^{1/3}-2}{x} APPLY L'HOSPITAL RULE.. \lim_{x \rightarrow 0} \frac{1/3(x+8)^{-2/3}}{1} =1/12 \]

OpenStudy (anonymous):

until now I had never heard of l'hostpital's rule. But I just wiki'd it. Interesting. Thank you guys! Much appreciated :)